suppose you take a sheet of paper of the ordinary variety (about 0.1mm or .004inches thick) and fold it in half creating paper which is about 0.2mm or .008 in thick- about the thickness of your finger nail. Using the exponetial growth formula and converting all answers to feet, determine the thickness of the paper after folding
14 times?
17 times?
22 times?
how would i put this into and exponential growth chart?

Answer by stanbon(75887)

14 times: 2^14(0.004)=65.536mm
Formul: 1mm=(1/1000)meters
65.536mm=(1/1000)65.536mm=0.065536 meters
Formula: 1 meter=39.37 inches
0.065536 meters = 39.37(0.065536)=2.58 inches
Formula: 1 inch = 1/12 ft.
2.58 inches = 2.58(1/12) ft.= 0.215 ft

Putting that all together:
thickness in feet for x folds=[(0.004)(39.37)]/[1000(12)]]2^x
Simplifying you get:
thickness in feet for x folds=(1.31233333333..E-5)2^x
Using that formula for 14 folds you get 0.21501 ft
17 folds gives 1.7201 ft.
22 folds gives 55.043 ft.
50 folds gives 1.5E10 ft=2,840,909.091 MILES
55 folds will get you to the sun
56 folds will get you to the sun and back to the earth.
Cheers,
Stan H.